OpenStudy (anonymous):

find the quotient. b^2+2b over 7 divided by b+2 over b plzzzzzzzz help!

OpenStudy (anonymous):

\[\frac{b^2+2b}{7}\times \frac{b}{b+2}\] is a start

OpenStudy (anonymous):

then factor the numerator and cancel \[\frac{b(b+2)\times b}{7(b+2)}=\frac{b^2}{7}\]

OpenStudy (anonymous):

so the awnser is b^2 over 7????